OpenStudy (anonymous):

what is the equation in standard form of the line which passes through -2,6 and has a slope of -1

OpenStudy (jackellyn):

y-6=-1(x-(-2)) y-6=-1(x+2) y-6=-x-2 y=-x-2+6 y=-x+4 @ozaze